# Competing Effects of Local Solvation Structures on Chemical Shift Changes of Liquid Electrolyte

### v1

@@ Abstract
Understanding the solvation structure of electrolytes is critical for optimizing the electrochemical performance of rechargeable batteries, as it directly influences properties such as ionic conductivity, viscosity, and electrochemical stability. The highly complex structures and strong interactions in high-concentration electrolytes make accurate modeling and interpretation of their “structure-property” relationships even more challenging with spectroscopic methods. In this study, we present a machine learning-based approach to predict dynamic 7Li NMR chemical shifts in LiFSI/DME electrolyte solutions. Additionally, we provide a comprehensive structural analysis to interpret the observed chemical shift behavior in our experiments, particularly the abrupt changes in 7Li chemical shifts at high concentrations. Using advanced modeling techniques, we quantitatively establish the relationship between molecular structure and NMR spectra, offering critical insights into solvation structure assignments. Our findings reveal the coexistence of two competing local solvation structures that shift in dominance as electrolyte concentration approaches the concentrated limit, leading to anomalous reverse of 7Li NMR chemical shift in our experiment. This work provides a detailed molecular-level understanding of the intricate solvation structures probed by NMR spectroscopy, leading the way for enhanced electrolyte design.
@@ 1 Introduction
Electrolytes facilitate the transfer of ions between the anode and cathode, and their solvation structure can influence various electrochemical properties, such as conductivity,1, 2 viscosity,3 and battery performance, including the electrochemical stable potential window (ESPW),4, 3 coulombic efficiency (CE)5, and cycling reversibility.6, 7 The relative strength of interactions between the cation–solvent and cation-anion can give rise to different types of solvation structures in the electrolytes.8, 9, 10, 11, 12, 13, 14 These structures significantly affect the formation of electrode-electrolyte interface,6, 15, 16, 17, 18 the desolvation and diffusion process,19, 20 and reaction mechanism.3, 21, 22 In recent years, highly concentrated electrolytes (HCEs) and localized highly concentrated electrolytes (LHCEs) are widely employed in electrochemical energy storage systems 9, 23, where the intricate solvation structures and strong interactions between solutes and solvents play a pivotal role in achieving high energy storage densities.8, 24, 25 These factors make the accurate modeling and interpretation of their ”structure-property” relationships particularly challenging. Consequently, revealing the solvation structures of electrolytes is the central topic in understanding these processes of interest.26, 27, 28
@@
In the experiments, there have been several characterization methods used to illustrate solvation structures, including vibrational and nuclear magnetic resonance (NMR) spectroscopies. Specifically, Raman spectroscopy has often been employed to reveal the intensity of cation-anion associations.29, 8, 30, 31, 13, 32, 33, 34, 9 NMR spectroscopy, a non-destructive and atom-specific technique, is particularly well-suited for investigating the molecular details of the chemical environments of the particular nucleus within solvation structures, since its signal is sensitive to the local chemical environment. 35, 36, 37, 38, 32, 39, 40, 41. Apart from that, the experimental NMR spectroscopy reveals much more information about correlation times,42, 41 relaxation times43, 44 and exchange dynamics of the chemical components.45 Nevertheless, how to connect the observed spectral changes to their intrinsic molecular structural changes is a very challenging task. For instance, experimental NMR spectroscopy has observed chemical shift variations with changing concentration, particularly in the context of ion coordination in glyme electrolytes.1, 46, 37 While its solvation structural changes are hard to connect to the observed chemical shift. Some calculations, which employ density functional theory (DFT) calculation can provide some limited information into these trends for electrolytes, often employing cluster extraction with some complicated sampling methods.47, 39, 36, 35, 40 There have been many static calculations attempting to explain the structure-spectrum relationship, using the Amsterdam Density Functional (ADF) package35, 37, 38, 36, 48, 39 or Gaussian package47, 40 to calculate 1H, 25Mg, 43Ca, 67Zn, 17O, 7Li NMR spectrum. However, experimental chemical shifts reflect a weighted average from various local sites, combining local structural and dynamic information. This averaging complicates signal resolution and makes it more challenging to evaluate the structure-spectrum relation.47, 49, 43
@@
Molecular dynamics (MD) simulations allow for tracking the dynamic structural changes in various types of electrolytes, with classical force field models,50, 47, 13 first-principles methods51, 52, 14 and machine learning potentials.53, 54 However, the direct connections between the molecular structures to the experimental spectral observables are still challenging, due to the heavy computational cost of obtaining a single spectral response from the obtained MD configuration. Moreover, the delicate choices of configurational sampling when utilizing the NMR DFT calculation are too complicated to be implemented in the complex electrolyte system.47 Nevertheless, there are some attempts, by training on structural descriptors of solid state structures and their corresponding NMR chemical shifts, machine learning models can enhance the speed of chemical shift predictions while maintaining high accuracy.55, 56, 49, 57, 58 Therefore, a robust computational method is essential, as a clear consensus on the relationship between the dynamic structural characteristics of electrolytes and their experimental spectral observables has not yet been established. Furthermore, validating the simulated structures and linking them to experimental observables remains a challenging task, one that could serve as a benchmark for the simulation.
@@
In this paper, we propose a novel approach to connect the microscopic molecular structures and the spectroscopic features of the electrolyte. First, we combine two machine learning (ML) models for calculating dynamic NMR shifts: one MLP model for accelerating configuration sampling and another Neural Network (NN) model for rapid chemical shift prediction. To elaborate, we use the MLP model to generate a range of concentration trajectories. Subsequently, we conduct extensive sampling of electrolytes through prolonged MLMD simulations, creating chemical shift datasets from infrequent configurations within these simulations using the DFT method. The NN model is then trained and validated with this DFT-generated data, where the input features are the structural descriptor, such as Local Many-body Tensor Representation (LMBTR),59, 60 and the outputs are the chemical shifts. Finally, we use the NN-NMR model to calculate the dynamic 7Li NMR spectra for various concentrations of LiFSI in DME solutions. We successfully reproduced the experimental observed anomalous reverse of 7Li NMR chemical shift in our experiment when electrolyte concentration approaches the concentrated limit. Furthermore, by introducing one local structural parameter, called local structural index (LSI),61, 62, 63 we quantitatively establish the structure-spectrum relationship, which helps clarify the solvation structure assignments. To this end, We find two different types of local solvation structures competing with each other, as electrolyte concentration approaches the concentrated limit, leading to anomalous reverse of 7Li NMR chemical shift in our experiment. As such, we established a comprehensive molecular-level understanding of the delicate microscopic molecular structures observed by using NMR spectroscopy. With our approach, one can track the complicated solvated structures changes in the electrolytes with different conditions, enabling precise control over the stability and solubility properties of electrolytes.
@@ 2.1 Theoretical NMR Calculation
The methodology is illustrated in Fig. 1. For the structural exploration section, Machine Learning Molecular Dynamics (MLMD) simulations are conducted under the NVT ensemble with various initial structures of different concentrations pre-equilibrated from classical MD simulations. The Nose-Hoover thermostat is employed for generating the NVT ensembles. The temperature is set to 300 K and each simulation lasts 10 nanoseconds (ns) with a 0.5 fs time step. More details about the pre-equilibrium and the MLMD simulations are given in Sect. I of the Supporting Information (SI). For the construction of the NN model, we sparsely sample configurations from MLMD trajectories. Then we extract the first solvation shells around the Li+ ions as clusters and label them, resulting in a DFT dataset containing about 28,000 7Li chemical shifts. All DFT calculations for NMR are performed by using Gaussian 16.64 NMR calculations for clusters are conducted at the revTPSS65, 66/pcSseg-167 level using the polarizable continuum model (PCM) solvation model.68
@@
The principle of cluster extraction is to capture the first solvation shell around the given Li+ ion while maintaining the integrity of surrounding molecules. The first solvation shell is identified by the first minimum position after the main peak in the radial distribution function (RDF) between Li and O shown in Fig. S3(c). We consequently extract complete molecules that have atoms located within this region. However, in some cases, other nearby Li+ ions occupy the space within the first coordination shell, as shown in the RDF of Li+-Li+ at 4M in Fig. S3(d), which exhibits a gentle peak around . Thus, we grasp the entire group of central atoms bonded to each other as one long-chain cluster. Although the cutoff distance between Li+ ions is sufficiently set within the range of the first shell, we extend the distance to to ensure the accurate local environment, which corresponds to the prominent peak position of RDF between Li+ ions at 1-3 M.
@@
For the construction of the NMR dataset, we use the LMBTR descriptor to encode the structure. The LMBTR vectors for the Li+ local environments are generated using the Dscribe package60, with a cutoff distance rcut of 6 Å, parameters of k2 and k3 are listed in Tab. S2 in SI. The NN prediction model is initialized and trained in PyTorch, utilizing a hidden layer with three sequential layers, each containing 256 nodes. The learning rate is initialized at 103 and decayed inversely with each epoch at a decay rate of 32. The model undergoes validation using a 9/1 split for the training and testing sets, and an early stopping mechanism is employed to prevent overfitting. We validate our NN-based NMR model using LiFSI/DME solutions. As shown in Fig. S5, the root-mean-square-error (RMSE) for the 7Li isotropic values in the test set is ppm. Details of the NN model are provided in Sect. V of SI. It’s worth mentioning that the NN model demonstrates outstanding efficiency, capable of rapidly predicting 7Li shifts for large datasets, which would be much more time-consuming and costly with traditional DFT methods. The prior limitation to the model’s prediction speed lies in the time required to convert molecular structures into descriptor vectors, which scales with the number of configurations and the dimension of vectors.
@@
After obtaining the NN models, we predict the NMR spectra based on MD simulation trajectories. As shown in Fig. 1(b), we generate four 10 ns MLMD trajectories for 1, 2, 3 and 4 M, respectively. We grasp the snapshots in regular intervals to ensure the number of Li+ ions is approximately 90,000 for every concentration, which is enough to sample the chemical spaces. These snapshots are then encoded by the LMBTR descriptor as inputs for the NN model to predict NMR chemical shifts, which are aggregated into histograms of NMR chemical shifts corresponding to the respective concentrations. The Full Width at Half Maximum (FWHM) obtained from Lorentzian fitting for experimental data are 0.008, 0.017, 0.020 and 0.019 for 1 M, 2 M, 3 M and 4 M, respectively. We use these values to fit the associated histograms into NMR spectra, ensuring that the mean values of the histograms and the peak position of NMR spectra align (Fig. 2(b)). The Lorentz function involved is: , where is the FWHM, is the peak position, and A is a scaling factor.
@@ 2.2 Experimental NMR Measurement
Lithium bis(fluorosulfonyl)imide (LiFSI) and 1,2-dimethoxyethane (DME) are obtained from Duoduo Chemical Reagent Co., Ltd., with both chemicals having a purity of 99.99%. All sample preparation procedures for four electrolytes with different salt concentrations (1, 2, 3, and 4 M LiFSI) are conducted within an argon-filled glovebox to maintain anhydrous conditions. 7Li Nuclear Magnetic Resonance (NMR) spectroscopy is employed to characterize LiFSI/DME electrolytes of varying concentrations. Measurements were performed using a Bruker Avance III HD 400 MHz NMR spectrometer at room temperature. We conduct two repeated NMR measurements for every single concentration. The experimental NMR chemical shifts data are shown in Fig. 2(a).
@@ 3.1 NMR Spectra of LiFSI/DME Electrolytes with Different Concentrations
It is generally believed that in the dilute limit case, the solvated Li+/FSI- ions are fully separated by the solution DME molecules in the LiFSI/DME electrolytes, called solvent-separated ions pairs (SSIP) situation. As the concentration of LiFSI increases, the coordination numbers of the oxygen atoms of DME molecules around the Li+ ions change. More complicated solvation structures, like contact ion pairs (CIPs), and ion aggregates (AGGs) are emerged.3 This, in turn, affects the electron localization function (ELF) of the Li+ ions,69, 70, 50 resulting in altered shielding effects around the Li+ ions and changes in the observed NMR chemical shifts. This mechanism underlies the variations seen in the experimental NMR spectra for LiFSI/DME at different concentrations, shown in the Fig. 2(a), our measurement is consistent with the previous reports37, 32. As one can see, in the low concentration (1-3 M), the NMR chemical shift tends to shift upfield with the increasing LiFSI concentration. However, at 4 M LiFSI concentration, which is close to the saturated limit (5.3 M) of LiFSI/DME solution, the peak is downfield shifted to ppm, when compared with the results obtained on the 3 M case of ppm. The sudden changes of the chemical shift obviously reflect the changes of the Li+ ion solvation structures at various LiFSI concentrations.
@@
To reveal the microscopic structural picture concerning the changes of the NMR chemical shift, we calculate the dynamic NMR spectra with the help of NN models, the data is presented in Fig. 2(b). Similarly to the experimental data, the theoretical NMR chemical shift is shifted upfield at low concentration, whereas it shifts downfield at 4 M. Notably, our theoretical NMR spectra are in excellent agreement with the experimental data, as we clearly show the NMR chemical shift changing point at 3 M concentration. The similar turning points observed in both experimental and theoretical NMR spectra suggest structural changes in the electrolytes as the LiFSI/DME solution approaches its saturation limit. Notably, at this concentration limit, the number of FSI- ions becomes comparable to that of the solvent molecules (DME), resembling a scenario referred to as ’water-in-salts’ (WiSs) 51, 71. In this high-concentration regime, Li+ ions are presumed to coordinate more frequently with FSI- ions than with DME molecules, in contrast to the low-concentration case. Since Li+ ions tend to bond with oxygen atoms in either FSI- ions or DME molecules, the differing interaction strengths of the Li–O bonds with FSI- or DME disrupt the delicate balance of electron density around the Li+ ions. This, in turn, affects the dynamic shielding effects and the NMR chemical shift. Nonetheless, the accurate reproduction of chemical shift variations in NMR due to changes in LiFSI concentration demonstrates the reliability of our NN models as effective tools for interpreting 7Li NMR experiments involving electrolytes.
3.2 The Local Structures of the LiFSI/DME Electrolytes

### v2

# Decoding the Competing Effects of Dynamic Solvation Structures on Nuclear Magnetic Resonance Chemical Shifts of Battery Electrolytes via Machine Learning

@@ Abstract
Understanding the solvation structure of electrolytes is critical for optimizing the electrochemical performance of rechargeable batteries, as it directly influences properties such as ionic conductivity, viscosity, and electrochemical stability. The highly complex structures and strong interactions in high-concentration electrolytes make accurate modeling and interpretation of their “structure-property” relationships even more challenging with spectroscopic methods. In this study, we present a machine learning-based approach to predict dynamic 7Li NMR chemical shifts in LiFSI/DME electrolyte solutions. Additionally, we provide a comprehensive structural analysis to interpret the observed chemical shift behavior in experiments, particularly the abrupt changes in 7Li chemical shifts at high concentrations. Using advanced modeling techniques, we quantitatively establish the relationship between molecular structure and NMR spectrum, offering critical insights into solvation structure assignments. Our findings reveal the coexistence of two competing local solvation structures that shift in dominance as electrolyte concentration approaches the concentrated limit, leading to an anomalous reverse of 7Li NMR chemical shift in the experiment. This work provides a detailed molecular-level understanding of the intricate solvation structures probed by NMR spectroscopy, leading the way for enhanced electrolyte design.
@@ 1 Introduction
Electrolytes facilitate the transfer of ions between the anode and cathode, and their solvation structure can influence various electrochemical properties, such as conductivity,1, 2 viscosity,3 and battery performance, including the electrochemical stable potential window (ESPW),4, 3 coulombic efficiency (CE)5, and cycling reversibility.6, 7 The relative strength of interactions between the cation–solvent and cation-anion can give rise to different types of solvation structures in the electrolytes.8, 9, 10, 11, 12, 13, 14 These structures significantly affect the formation of electrode-electrolyte interface,6, 15, 16, 17, 18 the desolvation and diffusion process,19, 20 and reaction mechanism.3, 21, 22 In recent years, highly concentrated electrolytes (HCEs) and localized highly concentrated electrolytes (LHCEs) are widely employed in electrochemical energy storage systems9, 23, where the intricate solvation structures and strong interactions between solutes and solvents play a pivotal role in achieving high energy storage densities.8, 24, 25 These factors make the accurate modeling and interpretation of their “structure-property” relationships particularly challenging. Consequently, revealing the solvation structures of electrolytes is the central topic in understanding these processes of interest.26, 27, 28
@@
In the experiments, there have been several characterization methods used to illustrate solvation structures, including vibrational and nuclear magnetic resonance (NMR) spectroscopies. Specifically, Raman spectroscopy has often been employed to reveal the intensity of cation-anion associations.29, 8, 30, 31, 13, 32, 33, 34, 9 NMR spectroscopy, a non-destructive and atom-specific technique, is particularly well-suited for investigating the molecular details of the chemical environments of the particular nucleus within solvation structures, since its signal is sensitive to the local chemical environment.35, 36, 37, 38, 32, 39, 40, 41. Apart from that, the experimental NMR spectroscopy reveals much more information about correlation times,42, 41 relaxation times43, 44 and exchange dynamics of the chemical components.45 Nevertheless, how to connect the observed spectral changes to their intrinsic molecular structural changes is a very challenging task. For instance, experimental NMR spectroscopy has observed chemical shift variations with changing concentration, particularly in the context of ion coordination in glyme electrolytes.1, 46, 37 While its solvation structural changes are hard to connect to the observed chemical shift. Some calculations, which employ density functional theory (DFT) calculation can provide some limited information into these trends for electrolytes, often employing cluster extraction with some complicated sampling methods.47, 39, 36, 35, 40 There have been many static calculations attempting to explain the structure-spectrum relationship, using first-principles methods to calculate 1H, 7Li, 17O, 25Mg, 43Ca, 67Zn NMR spectra.35, 37, 38, 36, 48, 39, 47, 40 However, experimental chemical shifts reflect a weighted average from various local sites, combining local structural and dynamic information. This statistical averaging complicates signal resolution and makes it more challenging to evaluate the structure-spectrum relation.47, 49, 43
@@
Molecular dynamics (MD) simulations allow for tracking the dynamic structural changes in various types of electrolytes, with classical force field models,50, 47, 13 first-principles methods51, 52, 14 and machine learning potentials.53, 54 However, the direct connections between the molecular structures to the experimental spectral observables are still challenging due to the high computational cost of obtaining a single spectral response from the obtained MD configuration. Moreover, the delicate choices of configurational sampling when utilizing the NMR DFT calculation are too complicated to implement in the complex electrolyte system.47 Nevertheless, there are some attempts, by training on structural descriptors of solid-state structures and their corresponding NMR chemical shifts, machine learning models can enhance the speed of chemical shift predictions while maintaining high accuracy.55, 56, 49, 57, 58 Therefore, a robust computational method is essential, as a clear consensus on the relationship between the dynamic structural characteristics of electrolytes and their experimental spectral observables has not yet been established. Furthermore, validating the simulated structures and linking them to experimental observables remains a challenging task, one that could serve as a benchmark for the simulation.
@@
In this paper, we propose a novel approach to connect the microscopic molecular structures and the spectroscopic features of the electrolyte. First, we combine two machine learning (ML) models for calculating dynamic NMR shifts: one MLP model for accelerating configuration sampling and another Neural Network (NN) model for rapid chemical shift prediction. To elaborate, we use the MLP model to generate a range of concentration trajectories. Subsequently, we conduct extensive sampling of electrolytes through prolonged MLMD simulations, creating chemical shift datasets from infrequent configurations within these simulations using the DFT method. The NN model is then trained and validated with this DFT-generated data, where the input features are the structural descriptor, such as Local Many-body Tensor Representation (LMBTR),59, 60 and the outputs are the chemical shifts. Finally, we use the NN-NMR model to calculate the dynamic 7Li NMR spectra for various concentrations of LiFSI in DME solutions. We successfully reproduce the anomalous reverse of 7Li NMR chemical shift observed in our NMR experiment when electrolyte concentration approaches the concentrated limit. Furthermore, by introducing one local structural parameter, called local structural index (LSI),61, 62, 63 we quantitatively establish the structure-spectrum relationship, which helps clarify the solvation structure assignments. To this end, we find two different types of local solvation structures competing with each other, as electrolyte concentration approaches the concentrated limit, leading to the anomalous reverse of 7Li NMR chemical shift in our experiment. As such, we establish a comprehensive molecular-level understanding of the delicate microscopic molecular structures observed by using NMR spectroscopy. With our approach, one can track the complicated changes of solvation structures in the electrolytes with different conditions, enabling precise control over the stability and solubility properties of electrolytes.
@@ 2.1 Theoretical NMR Calculation
The methodology is illustrated in Fig. 1. For the structural exploration section, Machine Learning Molecular Dynamics (MLMD) simulations are conducted under the NVT ensemble with various initial structures of different concentrations pre-equilibrated from classical MD simulations. The Nose-Hoover thermostat is employed to generate the NVT ensembles. The temperature is set to 300 K and each simulation lasts 10 nanoseconds (ns) with a 0.5 fs time step. More details about the pre-equilibrium and the MLMD simulations are given in Sect. I of the Supporting Information (SI). For the construction of the NN model, we sparsely sample configurations from MLMD trajectories. Then we extract the first solvation shells around the Li+ ions as clusters and label them, resulting in a DFT dataset containing about 28,000 7Li chemical shifts. All DFT calculations for NMR are performed by using Gaussian 16.66 NMR calculations for clusters are conducted at the revTPSS67, 68/pcSseg-169 level using the polarizable continuum model (PCM) solvation model.70
@@
The principle of cluster extraction is to capture the first solvation shell around the given Li+ ion while maintaining the integrity of surrounding molecules. The first solvation shell is identified by the first minimum position 3.0 Å after the main peak in the radial distribution function (RDF) between Li and O shown in Fig. S3(c). We consequently extract complete molecules that have atoms located within this region. However, in some cases, other nearby Li+ ions occupy the space within the first coordination shell, as shown in the RDF of Li+-Li+ at 4 M in Fig. S3(d), which exhibits a gentle peak around 3.0 Å. Thus, we grasp the entire group of central atoms bonded to each other as one long-chain cluster. Although the cutoff distance between Li+ ions is sufficiently set within the range of the first shell, we extend the distance to 6.3 Å to ensure the accurate local environment, which corresponds to the prominent peak position of RDF between Li+ ions at 1-3 M.
@@
For the construction of the NMR dataset, we use the LMBTR descriptor to encode the structure. The LMBTR vectors for the Li+ local environments are generated using the Dscribe package60, with a cutoff distance rcut of 6 Å, the parameters of k2 and k3 are listed in Tab. S2 in SI. The NN prediction model is initialized and trained in PyTorch71, utilizing the hidden layer with three fully connected layers, each containing 256 nodes. The learning rate is initialized at 10-3 with a total of 1000 epochs, using the Adam optimizer. The NMR dataset is divided into training, testing, and validation datasets in an 8:1:1 ratio, with an early stopping mechanism employed to prevent overfitting. We validate our NN-based NMR model using LiFSI/DME solutions. As shown in Fig. S5, the root-mean-square-error (RMSE) for the 7Li isotropic values in the testing dataset is ppm. Details of the NN model are provided in Sect. VI of SI.
@@
After obtaining the NN models, we predict the NMR spectra based on MD simulation trajectories. As shown in Fig. 1(b), we generate four 10 ns MLMD trajectories for 1, 2, 3 and 4 M, respectively. We grasp the snapshots in regular intervals to ensure the number of Li+ ions is approximately 90,000 for every concentration, which is enough to sample the chemical space. These snapshots are then encoded by the LMBTR descriptor as inputs for the NN model to predict NMR chemical shifts, which are aggregated into histograms of NMR chemical shifts corresponding to the respective concentrations. The Full Width at Half Maximum (FWHM) obtained from Lorentzian fitting for experimental data are 0.008, 0.017, 0.020 and 0.019 for 1 M, 2 M, 3 M and 4 M, respectively. We use these values to fit the associated histograms into NMR spectra, ensuring that the mean values of the histograms and the peak position of NMR spectra align (Fig. 2(b)). The Lorentz function involved is: , where is the FWHM, is the peak position, and A is a scaling factor.
@@ 2.2 Experimental NMR Measurement
Lithium bis(fluorosulfonyl)imide (LiFSI) and 1,2-dimethoxyethane (DME) are obtained from Duoduo Chemical Reagent Co., Ltd., with both chemicals having a purity of 99.99%. All sample preparation procedures for four electrolytes with different salt concentrations (1, 2, 3, and 4 M LiFSI) are conducted within an argon-filled glovebox to maintain anhydrous conditions. 7Li Nuclear Magnetic Resonance (NMR) spectroscopy is employed to characterize LiFSI/DME electrolytes of varying concentrations. Measurements were performed using a Bruker Avance III HD 400 MHz NMR spectrometer at room temperature. We conduct three repeated NMR measurements for every single concentration. The experimental NMR chemical shifts data are shown in Fig. 2(a).
@@ 3.1 NMR Spectra of LiFSI/DME Electrolytes with Different Concentrations
It is generally believed that in the dilute limit case, the solvated Li+/FSI- ions are fully separated by the solution DME molecules in the LiFSI/DME electrolytes, called solvent-separated ions pairs (SSIP) situation. As the concentration of LiFSI increases, the coordination numbers of the oxygen atoms of DME molecules around the Li+ ions change. More complicated solvation structures, like contact ion pairs (CIPs), and ion aggregates (AGGs) are emerged.3 This, in turn, affects the ELF of the Li+ ions,64, 65, 72, 50 resulting in altered shielding effects around the Li+ ions and changes in the observed NMR chemical shifts. This mechanism underlies the variations seen in the experimental NMR spectra for LiFSI/DME at different concentrations, as shown in Fig. 2(a), our measurement is consistent with the previous reports37, 32. As one can see, at the low concentrations (1-3 M), the NMR chemical shift tends to shift upfield with the increasing LiFSI concentration. However, at 4 M LiFSI concentration, which is close to the saturated limit (5.3 M) of LiFSI/DME solution, the peak is downfield shifted to ppm, when compared with the results obtained on the 3 M case of ppm. The sudden changes of the chemical shift obviously reflect the changes of the Li+ ion solvation structures at various LiFSI concentrations.
@@
To reveal the microscopic structural picture concerning the changes of the NMR chemical shift, we calculate the dynamic NMR spectra with the help of NN models, and the data is presented in Fig. 2(b). Similarly to the experimental data, the theoretical NMR chemical shift moves upfield at low concentration, whereas it shifts downfield at 4 M. Notably, our theoretical NMR spectra are in excellent agreement with the experimental data, as we clearly show the NMR chemical shift turning point at 3 M concentration. The similar turning points observed in both experimental and theoretical NMR spectra suggest structural changes in the electrolytes as the LiFSI/DME solution approaches its saturation limit. Notably, at this concentration limit, the number of FSI- ions becomes comparable to that of the solvent molecules (DME), resembling a scenario referred to as ‘water-in-salts’ (WiSs) 51, 73, 53. In this high-concentration regime, Li+ ions are presumed to coordinate more frequently with FSI- ions than with DME molecules, in contrast to the low-concentration case. Since Li+ ions tend to bond with oxygen atoms in either FSI- ions or DME molecules, the differing interaction strengths of the Li–O bonds with FSI- or DME disrupt the delicate balance of electron density around the Li+ ions. This, in turn, affects the dynamic shielding effects and the NMR chemical shift. Nonetheless, the accurate reproduction of chemical shift variations in NMR due to changes in LiFSI concentration demonstrates the reliability of our NN models as effective tools for interpreting 7Li NMR experiments involving electrolytes.
3.2 Local Structures of the LiFSI/DME Electrolytes